OpenStudy (anonymous):

Locate the absolute extrema of f(x)=x^3 - (9/2)x^2 on the interval [-1,4]

OpenStudy (anonymous):

absolute max at x = -1 and min at x = 3 ??????

OpenStudy (anonymous):

take the derivative, set it equal to zero and solve to find the possibilities. then check the endpoints as well

OpenStudy (anonymous):

\[f'(x)=3x^2-9x\] \[3x^2-9x=0\] \[3(x^2-3)=0\] \[x=\pm\sqrt{3}\] unless i made a mistake

OpenStudy (anonymous):

so only candiate in your interval is \(\sqrt{3}\) since \(-\sqrt{3}<-1\)

OpenStudy (anonymous):

*candidate, aka critical point

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

3x^2 - 9x =0 is 3x(x - 3) =0 ? so x=0, x=3

OpenStudy (anonymous):

right? or what did i do wrong?

OpenStudy (anonymous):

yikes i must be tired.

OpenStudy (anonymous):

its ok lol

OpenStudy (anonymous):

yes the critical point is at \(x=3\) so your last job is to compute \(f(-1), f(3), f(4)\) biggest is max, smallest is min